OpenStudy (secretsvt):

Write x^2 − 6x + 7 = 0 in the form (x − a)^2 = b, where a and b are integers. (x − 1)^2 = 4 (x − 4)^2 = 3 (x − 3)^2 = 2 (x − 2)^2 = 1

OpenStudy (mathstudent55):

You need to complete the square. First, subtract 7 from both sides. What do you get?

OpenStudy (secretsvt):

x^2 -6x=-7

TheSmartOne (thesmartone):

Now a^2 + 2ab + b^2 = (a + b)^2 So, we have to find out what the b term is so that way we can add b^2 on both sides -6x = 2ab We know that a = x so -6x = 2xb What is b?

OpenStudy (mathstudent55):

Good. To complete the square, you take half of the coefficient of the x-term, and you square it. Then you add that to both sides. The x-term is -6x. What is half of -6?

OpenStudy (secretsvt):

-3

OpenStudy (mathstudent55):

Good. Now square -3. What do you get?

OpenStudy (secretsvt):

-9 right ?

OpenStudy (mathstudent55):